Two buses leave towns 594 km apart at the same time and travel toward each other. One bus travels 12 km/h slower than the other. If they meet in 3 hours, what is the rate of each bus?
Answer by mananth(16949)

let the speed of one bus be x kmph
the other bus speed will be x-12 kmph.
..
Time travel by both = 3 hours
..
Distance = rate * time
distance traveled by first bus = 3x
distance traveled by second bus = 3(x-12)
..
3x+3(x-12) = 594
3x+3x-36=594
6x-36=594
6x=558
x= 558/6
x=93 kmph speed of first bus
speed of second bus = 93-12=81 kmph
..